) Let X be a binomial variate with n =100, p = .1.0 Find the approximate value of

P 10( ≤ X ≤12) using: (5) 

 (i) normal distribution 

 (ii)poisson distribution 

 [You may like to use the following values. 

 P(Z ≤ 67.0 ) = .0 7486, P(Z ≤ 33.0 ) = .0 6293, P(Z ≤ )0 = ]


1
Expert's answer
2021-04-29T16:57:08-0400

(i) normal distribution


 Previously find mean and standard deviation

"\bold{E}[X]=np=100\cdot0.1=10"

"\sigma[X]=\sqrt{np(1-p)}=\sqrt{100\cdot0.1\cdot0.9}=\sqrt{9}=3"


"P (10\leq X\leq12)="


"F\left(\dfrac{12-\bold{E}[X]}{\sigma[X]}\right)-F\left(\dfrac{10-\bold{E}[X]}{\sigma[X]}\right)="


"F\left(\dfrac{12-10}{3}\right)-F\left(\dfrac{10-10}{3}\right)="


"F\left(\dfrac{2}{3}\right)-F(0)\approx0.7486-0.5000=0.2486"


(ii) Poisson distribution 


Previously find parameter of Poisson distribution

"\lambda=np=100\cdot0.1=10"


Formula for Poisson probability is

"P(k)=\dfrac{\lambda^ke^{-\lambda}}{k!}"



"P (10\leq X\leq12)="


"P(0)+P(1)+\ldots+P(10)+P(11)+P(12)-\bigg(P(0)+P(1)+\ldots+P(10)\bigg)="


"P(11)+P(12)="


"\dfrac{10^{11}\cdot e^{-10}}{11!}+\dfrac{10^{12}\cdot e^{-10}}{12!}="


"\dfrac{10^{11}\cdot e^{-10}}{11!}+\dfrac{10^{11}\cdot 10\cdot e^{-10}}{11!\cdot12}="


"\dfrac{10^{11}\cdot e^{-10}}{11!}\left(1+\dfrac{10}{12}\right)="


"\dfrac{10^{11}\cdot e^{-10}}{11!}\left(1+\dfrac{5}{6}\right)="


"\dfrac{10^{11}\cdot e^{-10}}{11!}\cdot\dfrac{11}{6}\approx0.2085"